Circular No. 9092

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003ma A. Rest, Physics Department, Harvard University, on behalf of the

SuperMACHO collaboration, reports the discovery of a supernova (at “VR” magnitude 20.5) at z = 0.289 behind the Large Magellanic Cloud on images obtained on 2003 Dec. 13 with the Cerro Tololo 4-m Blanco telescope in the course of the SuperMACHO microlensing survey (Rest et al. 2005, Ap.J. 634, 1103). The object, designated SN 2003ma, faded by only 2.9 mag in I over the first 4.7 years following peak brightness (I = 19.7). It is located at α = 5^{h}31^{m}01^{s}.878, δ = −70^{o}04′15′′.89 (equinox 2000.0), which is 0′′.1 south of its host galaxy. Spectroscopy obtained with the Gemini South telescope (+ GMOS) on 2003 Dec. 22.3 and with the Magellan Clay telescope (+ MagE) on 2008 Dec. 28.3 shows 2003ma to be a very luminous type-IIn supernova, similar to SN 1988Z (Turatto et al. 1993, MNRAS 262, 128). The 2008 spectrum shows an intermediate-width Hα line with a FWHM of 5800 km/s. Details will be given in a paper by Rest et al. (2009, Ap.J., submitted).

COMET P/2009 U6 (LINEAR) K. Kadota, Ageo, Saitama-ken, Japan, reports pre-discovery CCD im-

ages of this comet (cf. IAUC 9090), obtained with a 0.25-m f/5 reflector on Oct. 19.80–19.81 UT, showing the object at total mag 16.8 with a 0′.4 coma and a tail 0′.7 long in p.a. 285^{o}. The following elliptical orbital elements (from 29 observations, Oct. 19–30) appear on MPC 67414:

T = 2009 Aug. 11.839 TT  ω = 311.^{o}110
e = 0.59878  Ω = 112.915 }2000.0
q = 1.51357 AU  i = 14.680
a = 3.77239 AU n^{o} = 0.134518 P = 7.33 years

COMETS C/2009 K13 AND C/2009 K14 (SOHO) Further to IAUC 9091, additional Kreutz sungrazers have been found

on SOHO website images. C/2009 K13 was slightly diffuse and small (mag ∼ 8). C/2009 K14 was very diffuse and very faint (mag ∼ 8–8.5).

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 K13 May 31.171 4^{h}31.^{m}0 +20^{o}05′ C2 BZ 2009-O18
C/2009 K14  31.746 4 31.4 +20 02 C2 BZ 2009-O18
2009 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green